I have read from the list that 103 421 is a prime number
Is there a simple way to find if a number is a prime number or a composite number?
I am very weak in maths.
Thank you for your attention.

Answer by stanbon(75887)

Is there a simple way to find if a number is a prime number or a composite number?
I am very weak in maths.
----------------------------
Let's say you want to determine if 149 is a prime number.
---
1st: Find the square root of 149:
sqrt(149) = 12.2...
---
2nd: See if 149 is evenly divisible by any prime less than 12.2
149/2 = 74.5 no
149/3 = 47.666 no
149/5 = 29.8 no
149/7 = 21.2.. no
149/11 = 13.54.. no
----
3rd: If none of these primes divides 149, 149 is a prime number.
----
Cheers,
Stan H.